the length of a rectangle is 3 meters longer than the width. if the area is 28 square meters, what are the rectanles dimensions?
Answer by checkley77(12569)

L=W+3
LW=28
(W+3)W=28
W^2+3W-28=0
(W+7)(W-4)=0
W-4=0
W=4 ANS
L=4+3
L=7 ANS.
PROOF:
7*4=28
28=28